OpenStudy (anonymous):

Find the values of c and d that make the following function continuous for all x: {5x if x<1 f(x)= {cx^2+d if 1or=2

OpenStudy (anonymous):

replace x by 1 in the first and second expressions, and set them equal then replace x by 2 in the second and third, and set them equal as well. then you will be able to solve for c and d

OpenStudy (anonymous):

so first equation for example would say \[5=c+d\] good from there?

OpenStudy (anonymous):

you'll get two eqns,\c+d=5 and 4c+d=6.solve it for c and d